OpenStudy (downhill):

Points A(-2, 4), B(1, 3), C(4, -1) and D form a parallelogram. What are the coordinates of D?

OpenStudy (jiteshmeghwal9):



OpenStudy (downhill):

A. (5,5) B. (0,0) C. (1,-2) D. (1,0) E. (3,4)

OpenStudy (eliesaab):

Let me explain to you why the answer is (1,0)

OpenStudy (eliesaab):

As a vector A-B =D-C (-3,1) =(-4+x, 1+y) Hence -4+x =-3 y+1= 1

OpenStudy (eliesaab):

So y=0 and x=1

OpenStudy (downhill):

Thank you so much, I actually did what the other person's diagram said, but regardless thank you so much :)